OpenStudy (anonymous):

Solve for x : e^x e^(x+1) = 1

OpenStudy (anonymous):

i think its xe^2x+1

OpenStudy (anonymous):

but i could be wrong.

OpenStudy (anonymous):

But how would I solve for x? I don't understand the solving for x when there are more than one x in the equation ha

OpenStudy (kainui):

@aylorterb You should use the fact that anything to the zero power is just 1, so e^0=1. That will let you solve maybe?

OpenStudy (anonymous):

@Kainui how did you get e^0? which e would that be?

OpenStudy (kainui):

Ahhh, I'm saying you can turn your equation from: \[\LARGE e^x e^{(x+1)} = 1\] into \[\LARGE e^x e^{(x+1)} =e^0\] since \[\LARGE e^0=1\] Sorry for the confusion.

OpenStudy (anonymous):

Okay! No need to be sorry! Thank you!

OpenStudy (solomonzelman):

yes just mentioning tha e^0 comes from: a=e^(ln a) is the rule, and you know that ln(1)=0 So, 1=e^(ln 1) = e^0 then equate the exponents, after the second line that Kainui wrote.

OpenStudy (anonymous):

So.. ln (e^x) ln (e^(x+1)) = ln 1 x (x+1) = ln 1?

OpenStudy (kainui):

Not quite, but you're on the right path. The rule for a product is \[\LARGE \ln(ab)=\ln(a)+\ln(b)\] So when you separate the two e^x and e^(x+1) terms you should add them, not multiply.

OpenStudy (anonymous):

okay so it would be x+ (x+1) = ln 1 ? And then I could subtract 1 from both sides, and then divide by 2?

OpenStudy (solomonzelman):

YES! basically: \(\large\color{slate}{ e^xe^{x+1}=1 }\) \(\large\color{slate}{ e^{x+x+1}=1 }\) \(\large\color{slate}{ e^{x+x+1}=e^{\ln(1)} }\) \(\large\color{slate}{ e^{2x+1}=e^{0} }\) \(\large\color{slate}{ 2x+1=0 }\) and then solve for \(\large\color{slate}{ x }\)...

OpenStudy (anonymous):

Thank you guys so much!

OpenStudy (solomonzelman):

yes, just a very good things to know: \(\Large\color{teal}{ a=e^{\ln(a)} }\) \(\LARGE\color{slate}{ ... }\) yw!